# Elastic Motion Policy: An Adaptive Dynamical System for Robust and Efficient One-Shot Imitation Learning

### v1

@@ I INTRODUCTION
In this paper, we present a stable-guaranteed full-pose motion policy learned from very few (1-2) demonstrations that can adapt to new object configurations without new data at 30Hz, allowing for online adaptation even when the relevant task objects are moving. EMP is not just a standalone method that imitates demonstrations but rather a flexible framework that can combine many capabilities, such as task-level constraints, obstacle avoidance, and learning multi-step demonstrations. Our contributions are:
@@ II RELATED WORK
A common approach for making an IL system more robust is by augmenting the training dataset. Methods following the DAgger paradigm [10, 11, 12, 13] require an online interactive expert, adding burden during training. To remedy workload, data augmentation techniques that inject noise during data collection have been proposed [14, 15, 16]. Using NeRF, [17] focuses on creating novel views for visuomotor imitation learning. These methods, however, are not tractable when the environments are changing online and data inefficient.
@@ IV-A SE3 LPV-DS Policy Morphing
Elastic-DS [3] allows morphing an existing LPV-DS motion policy by posing geometric constraints on the GMM parameters $\Theta_{\gamma}$. The key idea is to leverage the geometric relationship between neighboring Gaussians in the GMM by specifying joints $\beta$, approximately a middle point between two neighboring Gaussians in the GMM,
$\mathbb{\Sigma}_{t}=\left(\mathbb{\Sigma}_{1}^{-1}+\mathbb{\Sigma}_{2}^{-1}\right)^{-1}$ (7)
$\beta_{i,12}=\mathbb{\Sigma}_{t}\left(\mathbb{\Sigma}_{1}^{-1}\mu_{1}+\mathbb{\Sigma}_{2}^{-1}\mu_{2}\right).$ (8)
After that, by utilizing Laplacian Editing,
@@
where $L$ is the graph Laplacian matrix, $\Delta$ is the Laplacian coordinates. Homogeneous transformation constraints related to relevant objects in the demonstrations $O_{\text{start}}$ and $O_{\text{end}}$ can be specified at the endpoints $T_{0,1}$ and $T_{n-1,n}$ while the other joints will adjust “elastically” to preserve local geometric relationships. Hence the name Elastic-DS. Given the new GMM, the motion policy is re-estimated with Eq. 3. For more details on the Euclidean Elastic-DS, refer to [3]. Following we focus on extending Elastic-DS to the quaternion space.
After the GMM is fit to the trajectory data, either with PC-GMM [4] or DAMM [24], each Gaussian can be treated as links connected via joints. By manipulating those joint angles, the original demonstration can now adjust to different behaviors. As introduced in [5], the mean and covariance of quaternion mixture model are defined on the tangent plane $T\mathbb{S}^{3}$, which is a 3-dimensional subspace embedded in $\mathbb{R}^{4}$. Since all the quaternion GMM means, $\tilde{\mu}_{k}$, are expressed in terms of the attractor, we can form a basis centered at the attractor by defining the null space of the attractor vector in quaternion space, where the null space of a 4-dimensional vector is a 3-dimensional hyperplane embedded in $\mathbb{R}^{4}$:
@@ IV-B Real-Time Convex Lyapunov Function Optimization
We show an example of both feasible and violation cases in Fig. 4. A data point is feasible when the gradient $\nabla V$ makes an obtuse angle with the velocity vector $\dot{x}$, while the violation occurs when $\dot{x}^{T}\nabla V>0$ or the angle is acute.
Figure 4: Examples of feasible and violation cases of Lyapunov stability constraints, where the colored contours are the level sets of the candidate Lyapunov function, and the light grey curve represents a trajectory ending at the attractor $x^{*}$.
We then formulate the estimation of $\mathbb{P}$ as the following convex optimization problem:
@@
where $\epsilon>0$ is a small value to ensure numerical stability. To show the convexity of our proposed optimization problem, we note that $\{x_{i},\dot{x}_{i},x^{*}\}$ are all known values from the demonstration data, and therefore the term ${\dot{x}}_{i}^{T}\mathbf{P}({x}_{i}-x^{*})$ is linear in $\mathbf{P}$. The convexity of the ReLU function $f=\max(0,x)$ can be directly shown as follows:
$f(\lambda x+(1-\lambda)y)$ $=\max(0,\lambda x+(1-\lambda)y)$ (19)
$\leq\max(0,\lambda x)+\max(0,(1-\lambda)y)$
$=\lambda f(x)+(1-\lambda)f(y)$
Given that the composition of a convex function with an affine mapping preserves the convexity [30], we conclude that the objective function in Eq. 18 is convex.
@@ V-A3 Multi-Step Trajectory Decomposition
We divide a long-horizon demonstration into multiple sub-goal trajectories to learn individual goal-oriented motions. Then, we stitch multiple motion policies with a one-hot activation function to compose the multi-step task, $\dot{x}=\sum_{k}\,\sum_{j}\delta\left(\xi\right)f_{kj}\left(x\right),$ where $f_{kj}\left(x\right)$ represents each of the single stable motion policies, and $\delta(\xi)$ is a one-hot activation function that activates the next DS when the current one reaches its attractor. The outermost sum indexing by $k$ is the decomposition at the task subgoal level. Preliminary results have been shown in our prior Elastic-DS paper [3] with manually defined segments. In this work, we utilize pre-trained visual representation models, Universal Visual Decompose (UVD)[35], for the decomposition. This approach fits well with EMP as it can decompose long-horizon tasks into goal-conditioned subtasks. By applying UVD directly on the recorded demonstration video, it decomposes the trajectory, and each segment is then processed individually following the data processing pipeline. In this case, different segments could have different keypose for the same object.
@@ V-B Runtime Policy Inference
After collecting and processing the demonstration trajectory we train the nominal policy using the SE3-LPVDS framework. By tracking the relevant object poses $O_{obj}$, we can retrieve the keypose in the robot base frame $O_{key,r}$ and use it as a geometric constraint in Eq. 9 to update the EMP policy online. Its output $\{\dot{x},w\}$ will be sent to a passive impedance controller [36], translating the desired pose velocity to joint torque for robot execution. For multi-step tasks, once reaching an attractor in a subtask, the one hot activation function $\delta(\xi)$ will activate the next subtask. We assume that the gripper will change to an open/closed state only at the end of each subtask.
@@ VI-B1 Book Placing
For this task, we provide a single demonstration as shown in Fig. 1, where the human demonstrator holds the book approaching the bookrack from the front opening and places it into the rightmost slot. Prior to execution, both the baseline and our approach learn the policy from the single nominal demonstration we provide. Both approaches then adapt their learned policy based on the scene changes defined by the keypose. We measure the success rate of the task completion over 10 runs each in both ID and OOD scenarios. In Table. II, we show that both methods succeed in the task when starting and ending are the same as the demonstration, i.e., ID scenarios. However, when we change the bookrack configurations more randomly and aggressively, the success rate of the baseline drops by half, while our approach manages to complete in more challenging OOD scenarios. We note that the specific constraints of this task come from both the placement pose and the approaching pose. For example, if the approaching pose is off, the book will collide with the bookrack and fail the task. By simply transforming (e.g. rotating and translating) a learned policy as in the baseline method, there is no guarantee that the task constraints are still valid in the transformed policy. On the other hand, EMP ensures that task constraints comply with the updated policy as enforced by Eq. 9, leading to higher success rate. An example of adaptation is shown in Fig. 1.
@@ VI-C Case Study: Book Placing with Obstacle Avoidance
We demonstrate the additional feature of EMP to perform obstacle avoidance in real time, ensuring safe execution in a dynamic environment. We reuse the same experiment setup as in the Book Placing task, and additionally place an mustard bottle as the obstacle to obstruct the path of booking placing motion. For simplicity, we model the obstacle as a sphere object with a $0.18m$ diameter. Given that the EMP follows the dynamical systems paradigm, we can easily morph the learned velocity field $\dot{x}=f_{p}(x)$ and obtain a new one $\dot{x}_{new}$ that incorporates the obstacle avoidance behavior by adopting the modulation approach from [37, 2]:
$\dot{x}_{new}=\mathbf{M}(x)f_{p}(x),$ (20)
where the modulation matrix $\mathbf{M}$ is constructed through eigenvalue decomposition with the normal and tangent directions of the obstacle boundaries. As shown in Fig. 10, we demonstrate that EMP’s compatibility with obstacle avoidance allows for more flexible behavior in real world tasks.
VII CONCLUSIONS
In this work, we present EMP, a full-pose one-shot imitation learning approach with stability guarantees that can adapt to novel environment configurations in real-time. Through various experiments, we show how EMP can be naturally combined with other components, such as obstacle avoidance and multi-segment, to empower DS-based imitation learning. One limitation is that our approach relies on accurate object tracking. In the future, we would like to explore how to use EMP in higher dimensional systems such as vision and multi-modal input. Additionally, we plan to learn constraint-aware policies such as conditioning on robot kinematics limits.
@@ VII CONCLUSIONS
where the new vector is the estimated desired orientation in quaternion space. We lastly compute the estimated angular velocity given the current state and time difference,

### v2

@@ I INTRODUCTION
In this paper, we present a stable-guaranteed full-pose motion policy learned from very few (1-2) demonstrations that can adapt to new object configurations without new data at 30Hz, allowing for online adaptation even when the relevant task objects are moving. EMP is not just a standalone method that imitates demonstrations but rather a flexible framework that can combine many capabilities, such as task-level constraints, obstacle avoidance, and learning multi-step demonstrations. We assume 3D meshes of the objects of interest are available and can be used within an object tracking perception framework. Our contributions are:
@@ II RELATED WORK
A common approach for making an IL system more robust is by augmenting the training dataset. Methods following the DAgger paradigm [10, 11, 12] require an online interactive expert, adding burden during training. To remedy workload, data augmentation techniques that inject noise during data collection have been proposed [13, 14, 15]. These methods, however, are not tractable when the environments are changing online and data inefficient.
@@
One-shot imitation learning has recently gained attention, exploring techniques such as self-attention and contrastive learning [21], invariant region matching [22], and interaction warping [23]. In contrast, we propose EMP as an alternative, emphasizing computational efficiency, converging behaviors, and real-time adaptability.
@@ IV-A SE3 LPV-DS Policy Morphing
Elastic-DS [3] allows morphing an existing LPV-DS motion policy by posing geometric constraints on the GMM parameters $\Theta_{\gamma}$. The key idea is to leverage the geometric relationship between neighboring Gaussians in the GMM by specifying joints $\beta_{i,k,k+1}$, a point in the same space as the policy, which is approximately a middle point between two neighboring Gaussians in the GMM,
$\mathbb{\Sigma}_{t}=\left(\mathbb{\Sigma}_{k}^{-1}+\mathbb{\Sigma}_{k+1}^{-1}\right)^{-1}$ (7)
$\beta_{i,k,k+1}=\mathbb{\Sigma}_{t}\left(\mathbb{\Sigma}_{k}^{-1}\mu_{k}+\mathbb{\Sigma}_{k+1}^{-1}\mu_{k+1}\right).$ (8)
where $\Sigma_{t}$ is the covariance of a distribution obtained by two neighboring Gaussians $\Sigma_{k}$ and $\Sigma_{k+1}$ obtained from DAMM [25]. After that, by utilizing Laplacian Editing,
@@
where $L$ is the graph Laplacian matrix, $\Delta$ is the Laplacian coordinates. Homogeneous transformation constraints related to relevant objects in the demonstrations $O_{\text{start}}$ and $O_{\text{end}}$ can be specified at the endpoints $T_{0,1}$ and $T_{n-1,n}$ while the other joints will adjust “elastically” to preserve local geometric relationships. Hence the name Elastic-DS. Given the new GMM, the motion policy is re-estimated with Eq. 3. For more details on the Euclidean Elastic-DS, refer to [3].
Similarly, if we can extract the joints $\beta_{i,k,k+1}$ in orientation space, we can apply the same operation as in Eq. 9 to manipulate those joints and adjust the original demonstration to different orientation behaviors. As introduced in [5], the mean and covariance of quaternion mixture model are defined on the tangent plane $T\mathbb{S}^{3}$, which is a 3-dimensional subspace embedded in $\mathbb{R}^{4}$. Since all the quaternion GMM means, $\tilde{\mu}_{k}$, are expressed in terms of the attractor, we can form a basis centered at the attractor by defining the null space of the attractor vector in quaternion space, where the null space of a 4-dimensional vector is a 3-dimensional hyperplane embedded in $\mathbb{R}^{4}$:
@@ IV-B Real-Time Convex Lyapunov Function Optimization
We then formulate the estimation of $\mathbb{P}$ as follows:
@@
where $\epsilon>0$ is a small value to ensure numerical stability. Given that ReLU is a convex function, Eq. 18 is a convex optimization problem [31] that can be solved via QPs.
@@ V-A3 Multi-Step Trajectory Decomposition
We divide a long-horizon demonstration into multiple sub-goal trajectories to learn individual goal-oriented motions. Then, we stitch multiple motion policies together to compose the multi-step task, $\dot{x}=\sum_{k}\,\sum_{j}\delta\left(\xi\right)f_{kj}\left(x\right),$ where $f_{kj}\left(x\right)$ represents each of the single stable motion policies, and $\delta(\xi)$ is a one-hot activation function that activates the next DS when the current one reaches its attractor. The outermost sum indexing by $k$ is the decomposition at the task subgoal level. Preliminary results have been shown in our prior Elastic-DS paper [3] with manually defined segments. In this work, we utilize pre-trained visual representation models, Universal Visual Decompose (UVD)[36], for the decomposition. This approach fits well with EMP as it can decompose long-horizon tasks into goal-conditioned subtasks. By applying UVD directly on the recorded demonstration video, it decomposes the trajectory, and each segment is then processed individually following the data processing pipeline. In this case, different segments could have different keypose for the same object.
@@ V-B Runtime Policy Inference
After processing the demonstration, we train the nominal SE3-LPVDS policy. By tracking object poses $O_{obj}$, we update the EMP policy online using the keypose $O_{key,r}$ as a geometric constraint in Eq. 9. The output $\dot{x},w$} is passed to a passive impedance controller [37], converting desired velocities to joint torques. For multi-step tasks, reaching a subtask attractor triggers the next subtask via the activation function $\delta(\xi)$, with gripper state changes occurring only at subtask completion.
@@ VI-B1 Book Placing
For this task, we provide a single demonstration as shown in Fig. 1, where the human demonstrator holds the book approaching the bookrack from the front opening and places it into the rightmost slot. Prior to execution, both the baseline method SE(3)-LPVDS and our approach learn the policy from the single nominal demonstration we provide. Both approaches then adapt their learned policy based on the scene changes defined by the keypose. We measure the success rate of the task completion over 10 runs each in both ID and OOD scenarios. In Table. II, we show that both methods succeed in the task when starting and ending are the same as the demonstration, i.e., ID scenarios. However, when we change the bookrack configurations more randomly and aggressively, the success rate of the baseline drops by half, while our approach manages to complete in more challenging OOD scenarios. We note that the specific constraints of this task come from both the placement pose and the approaching pose. For example, if the approaching pose is off, the book will collide with the bookrack and fail the task. By simply transforming (e.g. rotating and translating) a learned policy as in the baseline method, there is no guarantee that the task constraints are still valid in the transformed policy. On the other hand, EMP ensures that task constraints comply with the updated policy as enforced by Eq. 9, leading to higher success rate. An example of adaptation is shown in Fig. 1.
@@ VI-C Case Study: Book Placing with Obstacle Avoidance
We demonstrate the additional feature of EMP to perform obstacle avoidance in real time, ensuring safe execution in a dynamic environment. We reuse the same experiment setup as in the Book Placing task, and additionally place an mustard bottle as the obstacle to obstruct the path of booking placing motion. For simplicity, we model the obstacle as a sphere object with a $0.18m$ diameter. We can morph the learned velocity field by adopting the modulation approach [38]: $\dot{x}_{new}=\mathbf{M}(x)f(x)$, where $\dot{x}_{new}$ is the new velocity that incorporates the obstacle avoidance behavior and the modulation matrix $\mathbf{M}$ is constructed through eigenvalue decomposition with the normal and tangent directions of the obstacle boundaries. As shown in Fig. 9, we demonstrate that EMP’s compatibility with obstacle avoidance allows for more flexible behavior in real world tasks.
VII CONCLUSIONS & DISCUSSIONS
In this work, we present EMP, a full-pose one-shot imitation learning approach with stability guarantees that can adapt to novel environment configurations in real-time. Through various experiments, we show how EMP can be naturally combined with other components, such as obstacle avoidance and multi-segment, to empower DS-based imitation learning. There are a few limitations and opportunities worth further exploration: (1) Our approach relies on accurate object tracking. In the future, we would like to explore using vision input directly to EMP (2). While the experiments specifically show capability in learning from a single demonstration, extending to multiple demonstrations is possible. By using DAMM [25], we can fit the multi-demonstration with a single target pose as a directed graph. The general Laplacian Editing can be used for graphs, so it would be suitable for morphing such a structure as well. To accommodate multi-demonstration with multiple goals, where different locally specific motions are needed for different parts of the objects, the learned GMM can adapt based on the local geometric constraints. However, information about the object shape is then needed. (3) In our evaluations, we did not observe significant distortions using a single tangent space to approximate the global quaternion manifold structure, as highlighted by [39]. This stability may partly be attributed to the conservative nature of the quaternion DS constrained by a QLF. Future work will explore Lie group formulations that naturally respect the manifold structure. Furthermore, while the P-QLF efficiently addresses a broad class of reaching motions, it inherently restricts more complex, highly non-linear motions. For these complex cases, future directions include exploring more expressive formulations such as Neural-ODE approaches [18] or leveraging contraction theory. The above limitations provide further exploration opportunities to enrich the EMP framework. We believe that EMP provides a way towards learning adaptive, robust, and time and data-efficient motion.
@@ VII CONCLUSIONS & DISCUSSIONS
where $(\hat{\mathbb{q}})^{des}$ is the estimated desired orientation in quaternion space. We then compute the desired angular velocity,